The ages of 100 persons are tabulated below. Find Q1, Q3, D2, D8, P20 and P60. Above what age


we have the oldest 10% persons?


Age(in years) 10-20 20-30 30-40 40-50 50-60 60-70 70-80


No. of persons 16 20 21 28 10 3 2


Note: the oldest 10% are above P90.


Ans.: P20 = 22, P60=41.071, the oldest 10% persons are of the age 55 and above.

1
Expert's answer
2021-12-10T12:45:13-0500

The frequency table for the data given above is as shown below.

class boundary f cf

10-20 16 16

20-30 20 36

30-40 21 57

40-50 28 85

50-60 10 95

60-70 3 98

70-80 2 100

Here, "n=100."

"i)"

"Q_1" is the "1^{st}" quartile defined by the formula,

"Q_1=l+({n\over4}-cf)\times({c\over f})" where,

"l" is the lower class boundary of the class containing "Q_1"

"cf" is the cumulative frequency of the class preceding the class containing "Q_1"

"c" is the width of the class in which "Q_1" lies

"f" is the frequency of the class containing "Q_1"

To determine the 1st quartile, we first find its position in order to find its class boundary as follows,

"Q_1=({n\over4})^{th} value=({100\over 4})^{th}=25^{th}value"

Therefore, the class boundary for the 1st quartile is 20-30

Thus,

"Q_1=20+(25-16)\times({10\over20})=20+4.5=24.5"


"ii)"

"Q_3" is the third quartile defined by the formula,

"Q_3=l+({3\over4}\times n-cf)\times({c\over f})" where,

"l" is the lower class boundary of the class containing "Q_3"

"cf" is the cumulative frequency of the class preceding the class containing "Q_3"

"c" is the width of the class containing "Q_3"

"f" is the frequency of the class containing "Q_3"

To determine the 3rd quartile, we first its position in order to find its class boundary as follows,

"Q_1=({3\over4}\times n)^{th} value=({3\over 4}\times 100)^{th}=75^{th}value"

Therefore, the class boundary for the 3rd quartile is 40-50

Thus,

"Q_3=40+(75-57)\times({10\over28})=40+6.4286=46.4286(4dp)"

Therefore, "Q_3=46.4286"


"iii)"

To find the 2nd decile; "D_2", we use the formula stated below,

"D_2=l+({(2\times n)\over10}-cf)\times({c\over f})" where,

"l" is the lower class boundary of the class containing "D_2"

"cf" is the cumulative frequency of the class preceding the class containing "D_2"

"c" is the width of the class in which "D_2" lies

"f" is the frequency of the class containing "D_2"

We first determine the class containing "D_2",

"D_2=({(2\times100)\over10})^{th} value=20^{th}value"

Therefore, the class where "D_2" can be found is 20-30

Hence,

"D_2=20+(20-16)\times({10\over20})=20+2=22"

Therefore, "D_2=22"


"iv)"

To find "D_8," we apply the formula below,

"D_8=l+({(8\times n)\over10}-cf)\times({c\over f})" where,

"l" is the lower class boundary of the class containing "D_8"

"cf" is the cumulative frequency of the class preceding the class containing "D_8"

"c" is the width of the class in which "D_8" lies

"f" is the frequency of the class containing "D_8"

The position of "D_8" is,

"D_8=({(8\times100)\over10})^{th}value.=80^{th}value"

The class in which "D_8" lies is 40-50

Thus,

"D_8=40+(80-57)\times({10\over28})=40+8.2143=48.2143(4dp)"

Therefore, "D_8=48.2143"


"v)"

To find the 20th percentile given as "P_{20}", we apply the following formula,

"P_{20}=l+({(20\times n)\over 100}-cf)\times({c\over f})" where,

"l" is the lower class boundary of the class containing "P_{20}"

"cf" is the cumulative frequency of the class preceding the class containing "P_{20}"

"c" is the width of the class in which "P_{20}" lies

"f" is the frequency of the class containing "P_{20}"

Let us determine the class where "P_{20}" lies by first finding its position,

"P_{20}={(20\times100)\over 100}=20^{th} position"

Now, "P_{20}" lies in the class 20-30

Thus,

"P_{20}=20+(20-16)\times({10\over 20})=20+2=22"

Therefore, "P_{20}=22"


"vi)"

To find the "60^{th}" percentile given as, "P_{60}", the following the formula is used,

"P_{60}=l+({(60\times n)\over 100}-cf)\times({c\over f})" where,

"l" is the lower class boundary of the class containing "P_{60}"

"cf" is the cumulative frequency of the class preceding the class containing "P_{60}"

"c" is the width of the class in which "P_{60}" lies

"f" is the frequency of the class containing "P_{60}"

To find the class where "P_{60}" lies, we first determine its position as follows,

"P_{60}={(60\times100)\over100}=60^{th}position"

Thus, "P_{60}" is in the class 40-50

Therefore,

"P_{60}=40+(60-57)\times({10\over28})=40+1.0714=41.0714"

Therefore, "P_{60}=41.0714(4dp)"


vii)

To find the age where the oldest 10% persons are above, we determine the 90th percentile given as, "P_{90}". We shall apply the formula below,

"P_{90}=l+({(90\times n)\over 100}-cf)\times({c\over f})" where,

"l" is the lower class boundary of the class containing "P_{90}"

"cf" is the cumulative frequency of the class preceding the class containing "P_{90}"

"c" is the width of the class containing "P_{90}"

"f" is the frequency of the class containing "P_{90}"

To find the class where "P_{90}" lies, we first determine its position as follows,

"P_{90}={(90\times100)\over100}=90^{th}position"

Thus, "P_{90}" is in the class 50-60

Therefore,

"P_{90}=50+(90-85)\times({10\over10})=50+5=55"

Hence, the oldest 10% persons are of the age 55 and above.